OpenStudy (anonymous):

x^3-3x^2-x+3 ... factor

OpenStudy (ajprincess):

what do you get when you factorise x^3-x?

OpenStudy (ajprincess):

@Laurenn_marie7

OpenStudy (anonymous):

x(x^2-1) ? @ajprincess

OpenStudy (ajprincess):

ya good:) Nw can u factorise -3x^2+3?

OpenStudy (anonymous):

3(x^2-1)

OpenStudy (ajprincess):

well small correction. it is -3(x^2-1)

OpenStudy (ajprincess):

Nw can u factorise x(x^2-1)-3(x^2-1)?

OpenStudy (anonymous):

that's the part i need help on!

OpenStudy (ajprincess):

think x^2-1 as y. Then u will have xy-3y Can u factorise this?

OpenStudy (anonymous):

so factor out x^2-1 or just -1?

OpenStudy (ajprincess):

factor out x^2-1

OpenStudy (anonymous):

x^2-1(x-3)

OpenStudy (ajprincess):

ya gud:) Nw x^2-1 can still be factoried. Can u do it? It is difference of squares.

OpenStudy (anonymous):

yes got it, thank you!! x-1 and x+1. thanks for all of the help! can you help me out with (2x-5)^4 ?

OpenStudy (anonymous):

@ajprincess

OpenStudy (ajprincess):

ya sure:) Well what do u want to do with it?

OpenStudy (anonymous):

factor it out! so it would be (2x-5)(2x-5)(2x-5)(2x-5) right?

OpenStudy (ajprincess):

ya

OpenStudy (anonymous):

and then I got (4x^2-10x-10x+25)(4x^2-10x-10x+25) and I'm stuck there

OpenStudy (ajprincess):

do u want to distribute it?

OpenStudy (anonymous):

Yes to just factor completely

OpenStudy (ajprincess):

factor completely will yield jst (2x-5)(2x-5)(2x-5)(2x-5)

OpenStudy (ajprincess):

if u want to distribute use pascal's triangle

OpenStudy (anonymous):

what is that?

OpenStudy (anonymous):

here i dont know how to explain it ill give you the answer 192/125

OpenStudy (anonymous):

it a fraction

OpenStudy (anonymous):

no the answer is 16x^4-160x^3+600x^2+625 but I'm not sure how they got all of that

OpenStudy (anonymous):

i thought that to but i went with that one

OpenStudy (anonymous):

well how do you get that from (4x^2-20x+25)(4x^2-20x+25) ?

OpenStudy (anonymous):

you right i factored it wrong sorry

OpenStudy (ajprincess):

pascal, triangle  The numbers in the pascal's triangle are the coefficients of the terms that make up the expression that u want to distribute. If u have an expression with an exponent then there will be n+1 terms when u distribute it. In ur case u have (2x-5)^4 since the exponent is 4 there will be 5 terms. nw c1(2x)^4+c2(2x)^3*5+c3(2x)^2*5^2+c4(2x)*5^3+c55^4 C1,c2,c3,c4 and c5 are the coefficients. frm the pascal's triangle we can find that they c1=1, c2=4,c3=6,c4=4 and c5=1. so c1(2x)^4+c2(2x)^3*5+c3(2x)^2*5^2+c4(2x)*5^3+c55^4 becomes (2x)^4-4*(2x)^3*5+6*(2x)^2*5^2-4*(2x)*5^3+5^4 =2^4x^4-4*2^3x^3*5+6*2^2x^2*5^2-4*2x*5^3+5^4 =16x^4-160x^3+600x^2-1000x+625